The age of a father is 4 years more than three times of his son’s age. Three years ago, father age was five times of him son’s age. Find the present age of each of them
1
Expert's answer
2020-08-11T19:01:08-0400

Suppose that "x" is the age of the father and "y" is the age of the son. Then we have the following system of equations:

"\left\{\begin{matrix}
 x=3y+4 \\
 (x-3)=5(y-3)
\end{matrix}\right."

We substitute expression for "x" from the first equation into the second and receive:

"3y+1=5y-15".

It implies that "y=8" . Using the first equation of the system we get "x=28" .

Answer: x=28, y=8.